# A new framework for nonlinear Kalman filters

### v10

@@ 1 Introduction
In control theory, the Kalman filter (KF) is a state estimation algorithm designed for linear systems. In such a system, a set of linear state transition functions describes how the states evolve, and a set of linear measurement functions describes the measured values of certain linear combinations of states. It is assumed that both the state transition and measurement functions contain some additive noise. The KF optimally combines the system knowledge and measurements to minimize the expected value of the mean-squared error of the state estimation. Additionally, when the process noise, measurement noise, and state estimation errors all follow a multivariate normal (Gaussian) distribution, the state estimation given by the Kalman filter is also the maximum likelihood estimation [1]. Due to these desirable features, KF is often referred to as the optimal linear state estimator. Since its birth in about 1960, KF has achieved massive success in navigation [2], motion control [3], statistics [4], signal processing [5], etc.
@@
Both EKF and EKF2 require the calculation of the Jacobian matrices of the state transition functions and measurement functions, which may not always be accessible in real-life systems. In 1997, Simon et al. introduced another extension of KF to nonlinear systems that requires no direct calculation of the Jacobian matrix and avoids the accuracy loss caused by linearization [8]. In later literature, this extension is usually referred to as the UKF [9]. UKF approximates the distribution of the states after nonlinear transformations by sampling several points (called sigma points) around the estimated mean of the states. These points go through the same transformation as the mean, and a more accurate estimation of the states’ distribution after the transformation can be obtained with the help of these extra sample points. In UKF, the sigma points are selected deterministically based on each state’s standard deviation. This technique guarantees that the sigma points are evenly distributed around the central point and can provide a more accurate state estimation. Similar ideas have also been used in several other nonlinear KFs, including cubature Kalman filter (CKF) [10], quadrature Kalman filter [11], square-root quadrature Kalman filter [12], etc. All these filters use deterministic ways to find several additional points around the central point (the estimated states from the previous iteration) and use those points to approximate the first two moments of the states after the nonlinear transformation. All these methods have some links with EKF2. As pointed out by Gustafsson et al. in [13], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the same state estimation as the EKF2, although the covariance estimation can still be a bit different. A disadvantage of these nonlinear KFs is that the number of sampling points is usually twice the number of states. Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
@@
While different extensions of the KF use various ways to approximate the first two moments of the states after a nonlinear transformation, they all follow the same framework as in the linear KF. In this paper, this framework is referred to as “the conventional framework” or “the old framework”. In such a framework, the state estimations in each iteration can be divided into two steps: Predict and Update. In the “predict” step, the state and covariance estimations are predicted using the state transition function and estimates from the previous iteration. Then, in the “update” step, the measurements correct the state predictions with a feedback gain (also known as the Kalman gain), which is designed to minimize the trace of the states’ covariance matrix after the update. After the states are updated, the state covariance matrix is also updated, ending the present iteration. Note that the covariance matrix will have the smallest possible trace after the update since that’s how the Kalman gain is selected.
While such a framework can guarantee the optimality of the state estimation for linear systems, it often leads to overconfident covariance estimation and less accurate state estimation when applied to nonlinear systems. The primary reason for such a phenomenon is the assumption that the Kalman gain minimizes the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions. In fact, as we will show later, the “update” step can sometimes make the state estimation of the nonlinear system less accurate, so the assumption about the optimality of the “update” step can be extremely harmful, as it makes the algorithm become overconfident in an inaccurate estimation.
The problem described above will become increasingly significant as the measurement noise decreases, and the measurements play a more critical role in the state estimation [14]. In some circumstances, it can even lead to a highly undesirable and counterintuitive phenomenon: the actual estimation error increases as the measurement noise decreases. The primary cause of this phenomenon is that the actual estimation errors drop much slower than the estimated standard deviations, and such overconfidence prevents the filter from correcting itself in later iterations. As the development of technology makes all types of sensors more and more accurate, we can reasonably predict that this problem will become increasingly severe if we stick to the conventional framework for nonlinear KFs.
@@
Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence. Although it is possible to mitigate the problem by approximating the measurement function at higher orders (e.g., using EKF2 instead of EKF), these algorithms take a significantly longer computational time, and the problem will still, in general, exist. The lack of theoretical analysis for nonlinear KFs makes finding a general solution even harder. For those few existing theoretical analyses [13, 14, 15], the results are based on the assumption that the higher-ordered terms in the Taylor expansions of the measurement functions can be ignored. This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the “update” step.
@@
In this study, we proposed a new framework that solves the problem of overconfidence. The key idea of the new framework is to add a step that reevaluates the actual effect of the Kalman gain after the state update. With this additional step, the state covariance estimation can be much more accurate, and any unhelpful update can be withdrawn. The main contributions of this paper are:
@@
The rest of the paper is organized as follows. Section 2 conducts a rigorous theoretical analysis of the problem of the conventional framework for nonlinear KFs and introduces a new framework that solves the problem. Section 3 presents a comprehensive comparison between the two frameworks. Specifically, both frameworks are combined with four different types of nonlinear KFs and validated on five distinct applications. The conclusions are drawn in Section 4. The detailed simulation setup and the detailed algorithms we used for validation can be found in the appendix.
@@ 2.1 General formulation of nonlinear Kalman filters
Consider a system with the following dynamics:
$x_{k}=f(x_{k-1},u_{k-1})+w_{k-1}$ (1)
@@
where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow a zero-mean multivariate distribution with covariance $Q_{k}$. The estimations of the states made in the previous step are denoted as $\hat{x}_{k-1|k-1}$. Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow a zero-mean multivariate distribution with covariance $P_{k-1|k-1}$. Whichever nonlinear KF is used, the first step of each iteration is called “predict”, which predicts the states using (1) and the estimations made in the previous time step. The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k}$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{k|k-1}$.
@@
The second step of each iteration for any nonlinear KF is called “update”, which corrects the predicted states based on the measurements. Suppose that the measurements of the states are as follows:
$z_{k}=h(x_{k})+v_{k}$ (2)
where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow a zero-mean multivariate distribution with covariance $R_{k}$. Given (2) and the predicted states, the measurements can also be predicted using a particular method specific to the type of nonlinear KFs. The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{y,k}$. The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy,k}$. The differences between the predicted and actual measurements are called the measurement residuals, denoted as $\tilde{y}_{k}$. Namely:
$\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$ (3)
If we assume that the measurement noises are independent from the predicted states, then $\tilde{y}_{k}$ should have zero mean, and their covariance $S_{k}$ can be calculated by:
$S_{k}=P_{y,k}+R_{k}$ (4)
KFs use the following equation to update the state estimation:
$\hat{x}_{k|k}=\hat{x}_{k|k-1}+K\tilde{y}_{k}$ (5)
where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (35), the covariance of $\hat{x}_{k|k}$ can be calculated as [16]:
$P_{k|k}=P_{k|k-1}+KS_{k}K^{T}-P_{xy,k}K^{T}-KP_{xy,k}^{T}$ (6)
KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of tr$(P_{k|k})$ in (6), we find that the trace is minimized when:
$K_{op}=P_{xy,k}S_{k}^{-1}$ (7)
where $K_{op}$ is the theoretical optimal value of the Kalman gain. When (7) is satisfied, (6) becomes:
$P_{k|k,op}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (8)
Equations (58) illustrate the key idea of KFs, which is to update the states in a way that minimizes the trace of the state covariance matrix after the update. However, it should be noted that these equations cannot be directly applied to systems with nonlinear measurement functions. The primary reason is that $P_{y,k},S_{k}$, and $P_{xy,k}$ can only be calculated if the actual state and the distribution of the states are known. Since neither of these holds in practice, the actual values for these covariance matrices are also unknown. In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that can make the conventional framework for nonlinear KFs overconfident and inaccurate.
@@ 2.2 The conventional framework and its problem
While calculating $P_{y,k},S_{k}$, and $P_{xy,k}$ at the actual state is impossible, it is possible to approximate their values based on the predicted states. Different nonlinear filters (e.g. UKF, EKF2, CKF) take different approaches to approximating these covariance matrices. In the subsequent analysis, we denote these approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ since they are all estimated based on the prediction. Although the detailed process of calculating ${\color[rgb]{1,0,0}P_{y,k|k-1}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (7) and (8) with the approximated values. Namely, the Kalman gain is selected as:
${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ (9)
and the state covariance matrix is updated by:
${\color[rgb]{1,0,0}P_{k|k,est}}$ $=P_{k|k-1}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k-1}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (10)
$=P_{k|k-1}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$

### v11

@@ 1 Introduction
In control theory, the Kalman filter (KF) is a state estimation algorithm designed for linear systems. In such a system, a set of linear state transition functions describes how the states evolve, and a set of linear measurement functions describes the measured values of certain linear combinations of states. It is assumed that both the state transition and measurement functions contain some additive noise. Among all linear estimators, the KF optimally combines the system knowledge and measurements to minimize the expected value of the mean-squared error of the state estimation. Additionally, when the process noise, measurement noise, and state estimation errors all follow a multivariate normal (Gaussian) distribution, the state estimation given by the Kalman filter is also the maximum likelihood estimation [1]. Due to these desirable features, KF is often referred to as the optimal linear state estimator. Since its birth in about 1960, KF has achieved massive success in navigation [2], motion control [3], statistics [4], signal processing [5], etc.
@@
Classical EKF and EKF2 formulations are usually presented in a way that requires explicit Jacobian matrices of the state-transition and measurement functions, which may not always be accessible or convenient to compute in real-world systems. Although derivative-free variants of the EKF and EKF2 have been proposed [8, 9], a particularly influential derivative-free extension of the KF to nonlinear systems is the UKF. In 1997, Julier and Uhlmann introduced a derivative-free extension of the KF to nonlinear systems that avoids explicit Jacobian computations and mitigates the accuracy loss caused by first-order linearization [10]. In later literature, this extension is usually referred to as the UKF [11]. The UKF approximates the mean and covariance of the states after nonlinear transformations by propagating a set of deterministically chosen sigma points around the estimated state mean. These sigma points are constructed from the square root of the state covariance matrix so that their weighted mean and covariance match the prior mean and covariance. After applying the nonlinear transformation to each sigma point, the transformed mean and covariance are recovered as weighted averages of the transformed points, which typically yields a more accurate approximation of the first two moments than a first-order linearization. Similar ideas have also been used in several other nonlinear KFs, including cubature Kalman filter (CKF) [12], quadrature Kalman filter [13], square-root quadrature Kalman filter [14], etc. All these filters use deterministic ways to find several additional points around the central point (the estimated states from the previous iteration) and use those points to approximate the first two moments of the states after the nonlinear transformation. All these methods have some links with EKF2. As pointed out by Gustafsson et al. in [9], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the same state estimation as the EKF2, although the covariance estimation can still be a bit different. A disadvantage of these nonlinear KFs is that the number of sampling points is usually twice the number of states. Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
@@
While different extensions of the Kalman filter (KF) approximate the first- and second-order moments of the state after a nonlinear transformation in different ways, they all retain the same algorithmic structure as the linear KF. In this paper, we refer to this structure as the “conventional framework” or “old framework”. In each iteration, the estimation process consists of two stages: Predict and Update. In the “predict” step, the state mean and covariance are propagated through the process model using the estimates from the previous iteration. In the “update” step, the incoming measurement is used to correct the state prediction through a feedback term known as the Kalman gain. In the linear KF, this gain is analytically derived to minimize the trace of the a posteriori covariance matrix, and the corresponding covariance is then updated consistently with this optimal gain.
While this framework is the optimal linear state estimator for linear systems, applying it to nonlinear measurement functions often results in overconfident covariance estimates and degraded state accuracy. The fundamental issue is that the Kalman gain is computed as though it still minimizes the posterior covariance trace, even though this optimality no longer holds when the required moments are only approximated. As we will show later, the “update” step may even increase the estimation error in nonlinear systems. Consequently, the implicit assumption of optimality in the “update” step can be very harmful, because the filter becomes increasingly confident in estimates that are systematically biased or inaccurate.
The overconfidence phenomenon described above is well known in the literature. Graphical illustrations of EKF overconfidence have been provided in [15], and in the context of simultaneous localization and mapping (SLAM), researchers have repeatedly observed that nonlinear KFs tend to produce overconfident, rather than conservative, state estimates, especially after a long time [16, 17, 18]. Moreover, this overconfidence issue becomes more pronounced when the measurement noise variance is small [19, 20], because the linearization (or other approximation) error then constitutes a larger fraction of the total error budget, making the filter’s covariance underestimate more severe. In some cases, this leads to a highly undesirable and counterintuitive behavior: the actual estimation error increases as the measurement noise decreases. The underlying reason is that the actual estimation errors decrease much more slowly than the estimated standard deviations, so the filter becomes overconfident and is less able to correct itself in later iterations. As sensor technology continues to improve and measurement noise is further reduced across many applications, we can reasonably expect this problem to become even more severe if we continue to rely on the conventional framework for nonlinear KFs.
@@
Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence and improve its estimation accuracy. Although it is possible to mitigate the problem by approximating the measurement function at higher orders (e.g., using EKF2 instead of EKF), these algorithms take a significantly longer computational time, and the problem will still, in general, exist. The lack of theoretical analysis for nonlinear KFs makes finding a general solution even harder. For those few existing theoretical analyses [9, 19, 21], the results are based on the assumption that the higher-ordered terms in the Taylor expansions of the measurement functions can be ignored. This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the “update” step.
Apart from modifying the structure of the filter itself, another widely used strategy for mitigating filter overconfidence is covariance inflation [22, 23]. In this approach, the estimated state covariance is deliberately enlarged, and the associated inflation factor is adjusted to compensate for under-dispersion and errors caused by nonlinearity. While covariance inflation can alleviate overconfidence, its performance depends heavily on the choice of the inflation factor, whose value typically requires extensive fine-tuning. This tuning can be computationally costly and problem-dependent, thereby undermining the universality of the approach [24, 25].
@@
In this study, we propose a new framework that solves the problem of overconfidence. The key idea of the new framework is to add a step that reevaluates the actual effect of the Kalman gain after the state update. With this additional step, the state covariance estimation can be much more accurate, and any unhelpful update can be withdrawn. The main contributions of this paper are:
@@
The rest of the paper is organized as follows. Section 2 conducts a rigorous theoretical analysis of the problem of the conventional framework for nonlinear KFs and introduces a new framework that solves the problem. Section 3 presents a comprehensive comparison between the two frameworks. Specifically, both frameworks are combined with four different types of nonlinear KFs and validated on five distinct applications. The conclusions are drawn in Section 4.
@@ 2.1 General formulation of nonlinear Kalman filters
Consider a discrete-time system with nonlinear dynamics
$x_{k}=f(x_{k-1},u_{k-1})+w_{k-1},$ (1)
where $x_{k}$ denotes the (unknown) true states at time step $k$, $u_{k-1}$ is the known input applied between time steps k1k-1 and $k$, and $w_{k-1}$ is the process noise, assumed zero-mean with covariance $Q_{k-1}$, and independent of $x_{k-1}$ and past measurements. At time step k1k-1, the filter maintains an a posteriori state estimate $\hat{x}_{k-1|k-1}$ with associated error covariance $P_{k-1|k-1}$.
@@
For any nonlinear KF variant, the first step of each iteration is the “predict” step, which propagates $\hat{x}_{k-1|k-1}$ and $P_{k-1|k-1}$ through the process model (1) to obtain the a priori quantities $\hat{x}_{k|k-1}$ and $P_{k|k-1}$. The predicted state estimate at time step $k$ is denoted by $\hat{x}_{k|k-1}$, and the corresponding estimation error $x_{k}-\hat{x}_{k|k-1}$ is assumed to follow a zero-mean multivariate distribution with covariance $P_{k|k-1}$. Formally,
$\begin{cases}\mathbb{E}[x_{k}-\hat{x}_{k|k-1}|z_{1:k-1}]=0,\\ P_{k|k-1}=\mathrm{Var}(x_{k}-\hat{x}_{k|k-1}|z_{1:k-1}),\end{cases}$ (2)
where $z_{1:k-1}$ denotes the past measurements. In the subsequent analysis, we work conditionally on the observation history. In particular, for any fixed realization of $z_{1:k-1}$, $P_{k|k-1}$ is a deterministic function of that realization; likewise, $P_{k-1|k-1}$ is deterministic given $z_{1:k-1}$. Unless stated otherwise, we suppress the conditioning in the notation for all means, covariances, and cross-covariances (e.g., $P_{k|k-1}$, $P_{z,k}$, $P_{xz,k}$, and $S_{k}$, which will be formally defined later).
@@
The second step of each iteration for a nonlinear KF is called the “update”, which corrects the predicted state based on the new measurement. The measurement model is
$z_{k}=h(x_{k})+v_{k},$ (3)
where $z_{k}$ is the measurement at time step $k$, and $v_{k}$ is the measurement noise, assumed zero-mean with covariance $R_{k}$. Given the predicted state, the measurement can be predicted (using a specific nonlinear KF algorithm), and we denote this predicted measurement by $\hat{z}_{k|k-1}$. The corresponding (noise-free) prediction error $h(x_{k})-\hat{z}_{k|k-1}$ is assumed zero-mean with covariance $P_{z,k}$. The cross-covariance between the state prediction error and the measurement prediction error is denoted by $P_{xz,k}$.
The difference between the actual and predicted measurements is called the measurement residual (or innovation), denoted by
$\tilde{z}_{k}:=z_{k}-\hat{z}_{k|k-1}.$ (4)
Since both $v_{k}$ and $h(x_{k})-\hat{z}_{k|k-1}$ are zero-mean, the innovation $\tilde{z}_{k}$ is also zero-mean. If, in addition, the measurement noise $v_{k}$ is assumed to be independent of the (noise-free) prediction error $h(x_{k})-\hat{z}_{k|k-1}$, the covariance of the innovation is
$S_{k}=P_{z,k}+R_{k}.$ (5)
The KF update for the state estimate is
$\hat{x}_{k|k}=\hat{x}_{k|k-1}+K_{k}\tilde{z}_{k},$ (6)
where $\hat{x}_{k|k}$ is the updated state estimate and $K_{k}$ is the Kalman gain at time step $k$. Using (4)–(6), the a posteriori error covariance can be written as [26]
$P_{k|k}=P_{k|k-1}+K_{k}S_{k}K_{k}^{T}-P_{xz,k}K_{k}^{T}-K_{k}P_{xz,k}^{T}.$ (7)
The Kalman gain $K_{k}$ is chosen to minimize $\mathrm{tr}(P_{k|k})$. By differentiating $\mathrm{tr}(P_{k|k})$ in (7) with respect to $K_{k}$, we obtain the optimal gain
$K_{k,\text{op}}=P_{xz,k}S_{k}^{-1},$ (8)
where $K_{k,\mathrm{op}}$ denotes the optimal Kalman gain. Substituting (8) into (7) yields the well-known expression
$P_{k|k,\mathrm{op}}=P_{k|k-1}-K_{k,\mathrm{op}}S_{k}K_{k,\mathrm{op}}^{T}.$ (9)
Equations (6)–(9) summarize the key principle of KF: the state estimate is updated in such a way that the trace of the a posteriori error covariance is minimized. However, for systems with nonlinear measurement functions $h(\cdot)$, these expressions involve covariance terms $P_{k|k-1}$, $P_{z,k}$, $S_{k}$, and $P_{xz,k}$ that, in general, are defined as expectations with respect to the full distribution of the state and cannot be uniquely determined from the first two moments alone. Since nonlinear KFs propagate only the mean and covariance of the state, this information is generally insufficient to recover the true values of these matrices, so they must be approximated. In fact, as we will show next, inaccurate approximations regarding these matrices are the key factors that can make the conventional framework for nonlinear KFs overconfident and inaccurate.
Before concluding this subsection, we note that the formulation above extends directly to non-additive noise. Specifically, if the process noise enters the dynamics as $x_{k}=f(x_{k-1},u_{k-1},w_{k-1})$, we can define the prediction-time augmented variable $\chi_{k-1}:=[x_{k-1}^{T}\;w_{k-1}^{T}]^{T}$. Likewise, if the measurement noise is non-additive, $z_{k}=h(x_{k},v_{k})$, we can define the update-time augmented variable $\eta_{k}:=[x_{k}^{T}\;v_{k}^{T}]^{T}$. Since the first two moments of these augmented variables are determined by the first two moments of $x_{k-1}$, $w_{k-1}$, and $v_{k}$, the same algorithmic machinery (e.g., linearization or sigma-point propagation) applies to $\chi_{k-1}$ and $\eta_{k}$ to approximate the first two moments of $x_{k}$ and $z_{k}$. Importantly, $w_{k-1}$ and $v_{k}$ are not treated as persistent states; they are i.i.d. zero-mean draws at each step and are included in the augmentation only to capture their nonlinear effect on the moments.
@@ 2.2 The conventional framework and its problem
While calculating the actual values of $P_{k|k-1},P_{z,k}$, $S_{k}$, and $P_{xz,k}$ is generally impossible by just using the first two moments of the states, it is possible to make some approximations. Without loss of generality, this problem of approximation can be formulated as follows.
The general approximation problem in nonlinear Kalman filters: Given an estimated mean $\hat{X}\approx\mathbb{E}[X]$ and an estimated covariance ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{x,\mathrm{est}}}\approx\mathrm{Var}(X)$ of the state $X$, and a nonlinear mapping $Y=g(X)$, we seek to approximate $\mathbb{E}[Y]$, $\mathrm{Var}(Y)$, and $\mathrm{Cov}(X,Y)$. The solution can be generally written as
$\begin{cases}\hat{Y}:=\Phi_{e}\bigl(g;\hat{X},P_{x,\mathrm{est}}\bigr)\approx\mathbb{E}[Y],\\[4.0pt] {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{y,\mathrm{est}}}:=\Phi_{y}\bigl(g;\hat{X},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{x,\mathrm{est}}}\bigr)\approx\mathrm{Var}(Y),\\[4.0pt] {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xy,\mathrm{est}}}:=\Phi_{xy}\bigl(g;\hat{X},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{x,\mathrm{est}}}\bigr)\approx\mathrm{Cov}(X,Y),\end{cases}$ (10)
where $\Phi_{e}$, $\Phi_{y}$, and $\Phi_{xy}$ denote the approximation mappings determined by the chosen nonlinear KF (e.g., EKF, UKF, CKF). The nonlinear mapping $g$ may represent either the state transition function $f$ (prediction step) or the measurement function $h$ (update step), so this approximation problem is typically solved at least twice in each iteration of a nonlinear KF. Specifically, $P_{k|k-1},P_{z,k}$, $S_{k}$, and $P_{xz,k}$ can be respectively approximated by:
$\begin{cases}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}:=\Phi_{y}(f;\hat{x}_{k-1|k-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k-1|k-1,\text{est}}})+Q_{k-1},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}}:=\Phi_{y}(h;\hat{x}_{k|k-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}),\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}}+R_{k},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}:=\Phi_{xy}(h;\hat{x}_{k|k-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}).\end{cases}$ (11)
Note that the subscript k|k1k|k-1 means that the covariance matrices are all estimated based on the predicted states $\hat{x}_{k|k-1}$.
Although the corresponding $\Phi_{y},\Phi_{s}$, and $\Phi_{xy}$ differ, all existing types of nonlinear KFs directly replace $P_{k|k-1}$, $S_{k}$, and $P_{xz,k}$ in (8) and (9) with the approximated values in (11). Namely, the Kalman gain is selected as
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}},$ (12)
and the state covariance matrix is updated by
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{est}}}$ $={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}^{T}}$ (13)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}.$